# VideoAuteur: Towards Long Narrative Video Generation

### v1

@@ 1 Introduction
Video generation [18, 39, 5, 6, 17, 49, 60] has recently witnessed remarkable advancements with diffusion [32, 2, 19, 55] and auto-regressive models [22, 41, 42, 51]. A primary objective is to generate coherent video clips from text prompts. In addition, a text-to-video model is able to serve as a foundation that can be adapted for various downstream applications, such as image animation [8, 52], video editing [11, 4], video stylization [20], etc.
@@
With the maturity of generating high-fidelity short video clips, researchers begin setting their sights on the next north-star: creating videos capable of conveying a complete narrative which captures an account of events unfolding over time. The literature has extensively highlighted narratives. The importance of narratives has been extensively documented in the literature, emphasizing their fundamental role for mankind. For example, Bruner argues that narratives are essential tools for organizing experiences and memories [3]. Similarly, Harari indicates that the ability to share narratives—stories and abstract ideas—has been pivotal in human development, setting humans apart from other animals and leading to the creation of nations and economic systems in our society [15]. It is not unexpected that all the key challenges associated with short-video will resurface in the context of long narrative video generation (NVG), along with new, unprecedented challenges across multi-clips, including constructing semantic consistency within complex event sequences, preserving object/character identity across multiple scenes, and more.
@@
A notble challenge researchers face is the lack of high-quality video data for learning narratives in video generation. While our community has developed many video datasets, most are unsuitable for NVG due to several reasons. First, most videos are tagged with descriptions that are partially or irrelevant to NVG. Second, even for the relevant descriptions, the quality may be insufficient, because these descriptions may be either too coarse or lack detailed actions needed for NVG. Finally, the type of video itself is a factor; not all videos contain meaningful narratives suitable for learning. Learning from videos that have meaningful, complete, and unambiguous narratives is essential not only for training but for evaluating NVG progress and benchmarking different approaches.
Partly due to the data challenge, progress in narrative video generation has been sluggish, especially when compared to story generation through a series of images, such as comics, cartoon or illustrations [21, 14, 54, 30]. Some of these images, when treated as keyframes, may be further processed into video clips [62, 59, 28, 63, 16]. A drawback of the keyframe-based approach is that the narrative video creation process is scattered across different modules and fine-tuned in multiple steps, making it difficult to optimize the overall process. Very recently, Yang et al. [54] proposed a story generation method using a vision-language model (VLM) to generate both images and text. Despite the promising results, there has yet to be a comprehensive study on using VLM models for narrative video generation.
@@
This paper contributes to advancing research in narrative video generation in two ways. First, we curate and annotate a large-scale video dataset focused on the cooking domain. The samples in our dataset are structured with clear narrative flows, each composed of sequential actions and visual states. Our dataset consists of approximately 200,000 video clips, with an average duration of 9.5 seconds per clip. We chose cooking videos due to their complete and unambiguous narratives, which are more objective to annotate and evaluate consistently. To address video copyright concerns, we source videos from existing video datasets, YouCook2 [61] and HowTo100M [31]. Beyond the video pre-processing such as quality filtering and captioning, we conduct a caption-action matching mechanism to extract narrative clips, following the strict, step-by-step process inherent in cooking tasks.
@@
Furthermore, we present a general auto-regressive pipeline for long narrative video generation, comprising two main components: a long narrative director and a visual-conditioned video generation model. The long narrative director produces a coherent narrative flow by generating a sequence of visual embeddings or keyframes that represent the story’s logical progression, as illustrated in Figure 1.
@@
Extensive experiments on our dataset demonstrate the effectiveness of the proposed pipeline for long narrative video generation. To sum up, our contributions are as follows:
@@
  • We build CookGen, a large, structured dataset and a comprehensive data pipeline designed to benchmark long-form narrative video generation. We will open-source the data along with the necessary functionalities to support future long video generation research.
  • We propose VideoAuteur, an effective data-driven pipeline for automatic long video generation. We emperically explore the design and training of an interleaved image-text auto-regressive model for generating visual states and a visual-conditioned video generation model.
@@ 2 Related Works
Text-to-Image/Video Generation Text-to-image [36, 33, 7, 56, 23, 34, 48] and video generation [18, 39, 5, 6, 17, 49, 60] have made remarkable progress to generate high-fidelity video clip of 5-10 seconds. For example, latent design [36] has become mainstream, balancing effectiveness with efficiency. Building upon this design, diffusion-based models like DiT [32], Sora [2], and CogVideo [19, 55] leveraged larger datasets and explored refined architectures and loss functions to enhance performance. In contrast, auto-regressive models such as VideoPoet [22] and Emu series [41, 42, 51] sequentially predict image or video tokens. Instead, our work focuses on the model’s ability to generate long narrative videos beyond 5 seconds.
Figure 2: CookGen contains long narrative videos annotated with actions and captions. Each source video is cut into clips and matched with the labeled “actions”. We use refined pseudo labels from ASR for Howto100M videos and use manual annotations for Youcook2 videos. We use state-of-the-art VLMs (i.e. GPT-4o and a finetuned video captioner) to provide high-quality captions for all video clips.
Interleaved Image-Text Modeling Interleaved image-text generation [9, 43, 13, 53, 1, 13] has garnered attention as a compelling research area that merges visual and textual modalities to produce rich outputs. Earlier approaches [35, 26, 35, 40] primarily relied on large-scale image-text paired datasets [12, 37] but were often confined to single-modality tasks, such as captioning or text-to-image generation. With the emergence of large language models (LLMs) [45], various vision-language models (VLMs) [25, 29, 50] have ushered in a new era of unified representations, leveraging well-curated datasets for interleaved generation. However, most existing works focus on the one-time generation and do not address the coherence of generated content, which is the focus of our work.
Narrative Visual Generation The narrative visual generation lies in ensuring consistency across generated images. With this motivation, several works like SEED-Story [54], StoryDiffusion [62], MovieDreamer [59], LVD [27], VideoDirectorGPT [28], Vlogger [63], Animate-a-story [16], VideoTeris [44] predominantly employ conditional generation in diffusion or auto-regressive models. Our work closely relates to the very recent work [54], which generates multimodal, long-form coherent narratives based on user-provided images and texts as story beginnings. However, we aim to generate coherent stories through videos, presenting greater identity and motion preservation challenges.
3 Long Narrative Video Data
@@ 3.1 Overview
We source over 30,000 raw videos about from two existing video datasets: YouCook2 [61] and HowTo100M [31]. Each video is filtered and cropped with processing to remove obvious logos or watermarks. Table 2 provides detailed information about the dataset statistics, video and clip details, and the train/validation partitioning.
@@ 3.2 Annotation and Processing
Data Source # Vid. (train/val) # Clips Clip Len. # Clips / Vid.
YouCook2 1333 / 457 $\sim$10K 19.6s 7.7
HowTo100M (subset) 30039 / 933 $\sim$183K 9.5s 5.9
3.3 Evaluation: Generation and Understanding
High-quality captions are essential for narrative visual generation. To assess the quality of our annotations, we evaluate them from two perspectives: inverse generation (§3.3.1) and visual understanding through VLM experts (§3.3.2).
3.3.1 Inverse Video Generation
This evaluation is motivated by the understanding that high quality captions, when combined with ground truth keyframes, more effectively reconstruct the original videos. We evaluate the dataset’s ability to reconstruct original videos using the annotated captions, with and without conditioning with ground truth keyframes. For this evaluation, we assess the validation set ($\sim$5,000 video clips) We measure reconstruction quality using FVD [47]. The results, shown in Table 3, indicate that our captions capture sufficient semantic information, enabling effective representation of the original videos. When generating with ground-truth keyframes, the video quality is very high and closely aligned with the original videos, as shown by the low FVD score (116.3). Without keyframes, the captions alone still provide reasonable alignment. Examples of reconstructed videos are included in the supplementary materials.
Validation Set w/. GT keyframe W/o. GT keyframe
# Clips FVD FVD
5504 116.3 561.1
Table 3: Inverse video generation. Evaluation of caption quality through inverse video generation with and without keyframes. FVD scores reflect reasonable video reconstruction quality.
3.3.2 Semantic Consistency across VLM Experts
GPT-4o Evaluation Human Evaluation
Score (0-100) Qwen2-VL-72B Ours Qwen2-VL-72B Ours
98.0 95.2 79.3 82.0
Figure 3: Long Narrative Video Director. The video director, a Visual Language Model (VLM), takes a user query (e.g., “How to cook a tuna sandwich?”) and an initial image-text pair as input. It then generates captions, actions, and visual states step-by-step. Each video clip is created using either visual embeddings or a keyframe derived from these embeddings. This study focuses on key design choices, such as the visual latent space, regression loss for visual embeddings, and the regression task.
@@ 4 Method
Given the text input, the task of long narrative video generation aims at generating a coherent long video $\mathcal{Y}\in\mathbb{R}^{H\times W\times F}$ that aligns with the progression of the text input sequentially. The $H$, $W$, and $F$ are generated videos’ height, width, and frame numbers. To achieve this, we propose VideoAuteur, a pipeline that involves two main components: a long narrative video director and visual-conditioned video generation. The long narrative video director is used to generate a sequence of visual embeddings or keyframes that capture the narrative flow (§4.1), while the visual-conditioned video generation generate video clips based on these visual conditions (§4.2).
4.1 Long Narrative Video Director
As shown in Figure 3, the long narrative video director generates a sequence of visual embeddings (or keyframes) that capture the narrative flow. In the following subsections, we first conduct a analysis of two kinds of video directors based on interleaved generation (§4.1.1) and language-centric visual generation (§4.1.2).
4.1.1 Interleaved Image-Text Director
The interleaved image-text director creates a sequence where text tokens and visual embeddings are interleaved, integrating narrative and visual content tightly. Using an auto-regressive model, it predicts the next token based on the accumulated context of both text and images, maintaining narrative coherence and aligning visuals with the text.
@@ 4.1.1 Interleaved Image-Text Director
where $\alpha$ and $\beta$ balance the contributions of cosine similarity and mean squared error to regress both scale and direction.
@@
This layered conditioning ensures coherence across the sequence, aligning actions, language, and visuals.
4.1.2 Language-Centric Keyframe Director
The interleaved auto-regressive model can also act as a language-centric director when reduced to using only text-based guidance. In this case, keyframes are synthesized using a text-conditioned diffusion model with only captions. For each caption $\mathbf{c}_{t}$, the diffusion model $D_{\text{text}}$ produces the visual state $\mathbf{x}_{t}=D_{\text{text}}(\mathbf{c}_{t})$. This text-only approach benefits from straightforward integration with off-the-shelf text-conditioned diffusion models (e.g., FLUX-1 [24]) and hence enjoys high-fidelity image generation. However, this approach structures the narrative and maintains visual consistency without regressed embeddings. Therefore, it lacks nuanced transitions between keyframes compared to the interleaved model, which enhances coherence by directly incorporating visual embeddings into the generation process.
@@ 4.2 Visual-Conditioned Video Generation
Using the sequence of actions $\mathbf{a}_{t}$, captions $\mathbf{c}_{t}$, and visual states $\mathbf{z}_{t}$ generated by the narrative director, we condition a video generation model to produce coherent long narrative videos. Unlike the classic Image-to-Video (I2V) pipeline that uses an image as the starting frame, our approach leverages the regressed visual latents $\mathbf{z}_{t}$ as continuous conditions throughout the sequence (see §4.2.1). Furthermore, we improve the robustness and quality of the generated videos by adapting the model to handle noisy visual embeddings, since the regressed visual latents may not be perfect due to regression errors (see §4.2.2).
@@ 4.2.1 Visual Conditions Beyond Keyframes
Traditional visual-conditioned video generation typically uses initial keyframes to guide the model, where each frame $\mathbf{x}_{t}$ is generated as $\mathbf{x}_{t}=D_{\text{visual}}(I_{t})$. Our interleaved auto-regressive director extends this by generating visual states $\mathbf{z}_{t}$ in a semantically aligned latent space, allowing direct conditioning without biases from a pretrained visual decoder, as shown in Figure 4. By using these regressed visual latents $\mathbf{z}_{t}$ directly, each frame is generated as $\mathbf{x}_{t}=D_{\text{visual}}(\mathbf{z}_{t})$, ensuring that the video accurately follows the narrative and enhancing consistency by relying on narrative-aligned embeddings rather than potentially biased keyframes (illustrated in Figure 5).
Figure 4: Visual-Conditioned video generation. Our interleaved auto-regressive director generates both text and visual conditions, enabling the video generation process to be conditioned either on keyframes (VAE embeddings) or on CLIP latents regressed by the interleaved director. We apply Gaussain noise, random masking and random shuffling as regularization during the training process to improve robustness with the imperfect visual embeddings.

### v2

@@ 1 Introduction
Video generation [20, 41, 5, 6, 19, 51, 62] has witnessed remarkable advancements with diffusion [34, 2, 21, 57] and auto-regressive models [25, 43, 44, 53]. A primary objective is to generate video clips from text prompts and supports various downstream applications, such as image animation [8, 54], video editing [11, 4], video stylization [23].
@@
With the maturity of generating high-fidelity short video clips, researchers begin setting their sights on the next north-star: creating videos capable of conveying a complete narrative which captures an account of events unfolding over time. The importance of narratives has been highlighted in the literature. For example, Bruner argues that narratives are essential tools for organizing experiences and memories [3]. The book Sapiens: A Brief History of Humankind emphasizes that the ability to share narratives (stories) has been pivotal in human development, setting humans apart from other animals [15].
@@
Long Narrative Video Generation (NVG) introduces several challenges. One particularly challenge is the scarcity of video data suitable for learning coherent narratives in video. While our community has developed many video datasets, most are unsuitable for NVG. First, most videos are tagged with descriptions that are partially to NVG. Second, even for the relevant descriptions, these descriptions may be either too coarse or lack detailed actions needed for NVG. Finally, not all videos contain meaningful narratives suitable for learning and can be well evaluated.
Consequently, video data with clear, complete, and meaningful narratives is crucial not only for training but also for evaluating and comparing NVG methods. However, compared to story generation through a sequence of images [24, 14, 56, 32], progress in narrative video generation has been relatively slow, partly due to the absence of standardized training and evaluation benchmarks.
@@
This paper contributes to advancing research in narrative video generation in two ways. First, we curate and annotate a large-scale video dataset on the cooking domain. The samples in our dataset are structured with clear narrative flows, each composed of sequential actions and visual states. Our dataset consists of approximately 200,000 video clips, with an average duration of 9.5 seconds per clip. We select cooking videos for their well-defined and less ambiguous narratives, making them more objective to evaluate consistently. To address video copyright concerns, we source videos from existing video datasets, YouCook2 [63] and HowTo100M [33]. We design various mechanisms to ensure high-quality videos and captions, organized in a structured storyboard format, as illustrated in Figure 1.
@@
Additionally, we propose a new auto-regressive pipeline for long narrative video generation, comprising three main components: a long narrative director, a rolling-context conditioned keyframe renderer, and a visual-conditioned video generation model. The long narrative director produces a coherent narrative flow by generating a sequence of visual embeddings or keyframes that represent the story’s logical progression. Building upon this, the rolling-context conditioned keyframe renderer utilizes a rolling history of reference images as contextual conditioning to generate high-quality keyframes with consistency. Finally, the visual-conditioned video generation model produces video clips based on these visual conditions to do narrative.
@@
Extensive experiments on the large-scale collected dataset demonstrate the effectiveness of the proposed pipeline for long narrative video generation. To sum up, our contributions are as follows:
@@
  • We construct CookGen, a large-scale, structured dataset accompanied by an effective data pipeline to benchmark long-form narrative video generation. The dataset along with the necessary functionalities will be opensourced to advance future research in the area.
  • We propose VideoAuteur, a novel approach for automated long video generation. It effectively bridges interleaved auto-regressive multimodal LLMs with pretrained DiTs, employing a rolling context strategy for enhanced generation quality and visual consistency.
  • Extensive experimental results and ablation studies show that VideoAuteur achieves the state-of-the-art performance in long narrative video generation.
@@ 2 Related Works
Text-to-Image/Video Generation Text-to-image [38, 35, 7, 58, 26, 36, 50] and video generation [20, 41, 5, 6, 19, 51, 62] have made remarkable progress to generate high-fidelity video clip of 5-10 seconds. For example, latent design [38] has become mainstream, balancing effectiveness with efficiency. Building upon this design, diffusion-based models like DiT [34], Sora [2], and CogVideo [21, 57] leveraged larger datasets and explored refined architectures and loss functions to enhance performance. In contrast, auto-regressive models such as VideoPoet [25] and Emu series [43, 44, 53] sequentially predict image or video tokens. Instead, our work focuses on the model’s ability to generate long narrative videos beyond a few seconds.
Figure 2: CookGen contains long narrative videos annotated with actions and captions. Each source video is cut into clips and matched with the labeled “actions”. We use refined pseudo labels from ASR for Howto100M videos and use manual annotations for Youcook2 videos. We use state-of-the-art VLMs (i.e. GPT-4o and an expert captioner) to provide high-quality captions for all video clips.
Interleaved Image-Text Modeling Interleaved image-text generation [9, 45, 13, 55, 1, 13] has garnered attention as a compelling research area that merges visual and textual modalities to produce rich outputs. Earlier approaches [37, 29, 37, 42] primarily relied on large-scale image-text paired datasets [12, 39] but were often confined to single-modality tasks, such as captioning or text-to-image generation. With the emergence of large language models  [47], various vision-language models  [28, 31, 52] have stepped in a new era of unified representations, leveraging well-curated datasets for interleaved generation. However, most existing works focus on the one-time generation and do not address the coherence of generated content, which is our focus.
Narrative Visual Generation Existing narrative visual generation primarily focuses on addressing challenges related to semantic and visual consistency. Recent approaches such as Narrative Visual Generation, VideoDirectorGPT [30], Vlogger [65], Animate-a-story [16], VideoTeris [46], IC Lora [22], Vlogger [65], and Animate-a-story [16] employ various methods to enhance semantic coherence and visual continuity. Unlike most prior methods that mainly focus on consistent image generation [22, 56, 64], our target is generating coherent narrative videos. While some works make efforts to be language-centric using text as conditions for video generation [65, 54] or appending with keyframes [61], different from these work, we propose an integrated approach that leverages multi-modal large language models (LLMs) in conjunction with in-context diffusion transformer models to ensure global narrative coherence, subsequently conditioning the video generation model.
3 CookGen: a Long Narrative Video Dataset
@@ 3.1 Overview
We source over 30,000 raw videos about from two existing video datasets: YouCook2 [63] and HowTo100M [33]. Each video is filtered and cropped with processing to remove corruptions. Table 1 provides detailed information about the dataset statistics, video and clip details, and the train/val partitioning.  Appendix B provides more details.
@@
Data Source # Vid. (train/val) # Clips Clip Len. # Clips / Vid.
YouCook2 1333 / 457 $\sim$10K 19.6s 7.7
HowTo100M (subset) 30039 / 933 $\sim$183K 9.5s 5.9
@@ 3.2 Annotation and Processing
Annotation Quality Reverification.
High-quality captions are essential for narrative visual generation. To verify the quality of our annotations, we build an evaluation pipline of inverse generation and visual understanding through VLM experts, which are detailed in Appendix §C.1 and §C.2.
Figure 3: Long Narrative Visual Condition Generation. (a) Interleaved Auto-regressive Director: an auto-regressive vison-language model, takes a user query (e.g., “How to cook a tuna sandwich?”) and an initial image-text pair as input. It then generates actions, captions, and visual states (i.e., visual embeddings) step-by-step. (b) Rolling Context Conditioned Render: Apart from the semantics consistency through interleaved generation, we use a rolling of reference images as direct context conditions to further improve visual consistency with a diffusion transformer model. With them, a long narrative video can be created using these generated visual conditions (i.e., visual embeddings and/or keyframes derived from the interleaved director and the keyframe render with rolling context conditioning.)
[…]
@@ 4 Method
Given the text input, the task of long narrative video generation aims at generating a coherent long video $\mathcal{Y}\in\mathbb{R}^{H\times W\times F}$ that aligns with the progression of the text input sequentially. The $H$, $W$, and $F$ are generated videos’ height, width, and frame numbers. To achieve this, we propose VideoAuteur, which involves three main components: an interleaved long narrative video director, a rolling-context conditioned keyframe renderer, and a visual-conditioned video generation model. The long narrative video director creates a sequence of language states and visual embeddings to represent the narrative flow (§4.1). A pretrained DiT model then renders keyframes using a rolling history of reference images as contextual conditioning (§4.2). Finally, the video generation model produces video clips based on these visual conditions (§4.3).
4.1 Long Narrative Interleaved Director
As shown in Figure 3, the long narrative video director generates a sequence of visual embeddings (or keyframes) that capture the narrative flow. The interleaved image-text director creates a sequence where text tokens and visual embeddings are interleaved, integrating narrative and visual content tightly. Using an auto-regressive model, it predicts the next token based on the accumulated context of both text and images. This helps maintain narrative coherence and align visuals with the text semantics.
@@ 4.1 Long Narrative Interleaved Director
where $\alpha$ and $\beta$ are hyper-parameters.
@@
This layered conditioning improves coherence across the sequence, aligning actions, language, and visuals.
4.2 Rolling Context Conditioned Render
While the interleaved auto-regressive director model can learn visual consistency, the CLIP representation space struggles to preserve fine visual details (e.g., character features, clothing patterns), as demonstrated in Appendix Figure 18. To address this limitation and improve generation quality, we employ a pretrained Text-to-Image diffusion transformer model to render high-quality keyframes, conditioning on a rolling history of reference images. The context length can vary dynamically from 1 to 3, balancing flexibility and efficiency when generating keyframes.
As illustrated in Figure 3, we use a rolling history of two reference images, $I_{t}$ and $I_{t-1}$. This setup is further conditioned by the tiled global caption
$\mathbf{c}_{\text{tiled}}=\text{tiled}(\mathbf{c}_{t-3},\,\mathbf{c}_{t-2},\,\mathbf{c}_{t-1},\,\mathbf{c}_{t}),$ (5)
the predicted visual embeddings $\mathbf{z}_{t}$ and $\mathbf{z}_{t-1}$, as well as the reference images $I_{t-3}$ and $I_{t-2}$.
$\text{D}(\mathbf{c}_{\text{tiled}},\,\mathbf{z}_{t-1},\,\mathbf{z}_{t},\,I_{t-3},\,I_{t-2})\rightarrow I_{t-1},I_{t},$ (6)
where $\text{D}(\cdot)$ denotes the diffusion model for synthesizing a new keyframe $I_{t}$ by integrating the rolling context of images, captions, and visual embeddings. This layered conditioning improves coherence across frames.
Flow Matching Loss. We employ a flow matching loss that aligns the learned drift function $f_{\theta}$ with the ground-truth path from $\mathbf{x}_{T}$ to $\mathbf{x}_{T+1}$. We define:
$\mathcal{L}_{\text{flow}}(\theta)\;=\;\mathbb{E}_{\mathbf{x}_{T},\,\mathbf{x}_{T+1},\,T}\Bigl{[}\bigl{\|}f_{\theta}(\mathbf{x}_{T},T)\;-\;\mathbf{v}(T)\bigr{\|}^{2}\Bigr{]},$ (7)
where $\mathbf{v}(T)$ denotes the ideal drift path that transitions $\mathbf{x}_{T}$ towards $\mathbf{x}_{T+1}$. This objective enforces consistency across frames without relying on a separate diffusion loss.
Figure 4: Visual-conditioned video generation. Our interleaved auto-regressive director and rolling context renderer generates both text and visual conditions, enabling the video generation process to be conditioned on keyframes (VAE embeddings) and CLIP latents. We apply Gaussian noise, random masking and random shuffling as regularization during the training process to improve robustness with the imperfect visual embeddings.
@@ 4.3 Visual-Conditioned Video Generation
Using the sequence of actions $\mathbf{a}_{t}$, captions $\mathbf{c}_{t}$,visual states $\mathbf{z}_{t}$ and keyframe ${I}_{t}$ generated by the interleaved director and rolling context conditioned render, we condition a video generation model to produce coherent long narrative videos. Unlike the classic Image-to-Video (I2V) pipeline that only uses an image as the starting frame, our approach leverages the regressed visual latents $\mathbf{z}_{t}$ as continuous conditions throughout the sequence (see §4.3.1). Furthermore, we improve the robustness and quality of the generated videos by adapting the model to handle noisy visual embeddings, since the regressed visual latents may not be perfect due to regression errors and keyframe uncertainty (see §4.3.2).
@@ 4.3.1 Visual Conditions Beyond Keyframes
Conventional visual-conditioned video generation typically uses initial keyframes to guide the model, where each frame $\mathbf{x}_{t}$ is generated as $\mathbf{x}_{t}=D_{\text{visual}}(I_{t})$. Our interleaved auto-regressive director supports generating visual states $\mathbf{z}_{t}$ in a semantically aligned latent space, allowing direct conditioning from a pretrained visual decoder, as shown in Figure 4. By using these regressed visual latents $\mathbf{z}_{t}$ directly, each frame is generated as $\mathbf{x}_{t}=D_{\text{visual}}(\mathbf{z}_{t})$. This follows the narrative and enhancing consistency by relying on narrative-aligned embeddings.